what is the probability of selecting a composite whole number from the whole numbers 2-15.
Answer by stanbon(48545)

what is the probability of selecting a composite whole number from the whole numbers 2-15.
---------
the composite numbers are 10,11,12,13,14,15
----------------------
the total number of numbers is 15-2+1 = 14
-----------------------
Prob(composite) = 6/14 = 3/7
=================================
Cheers,
Stan H.